OpenStudy (anonymous):

A local theatre sold 500 seats for an upcoming play. Ticket prices for adults are $52.00 and ticket prices for children are $22.00. If the theatre made $20,690.00 for the show, how many adult and children attended?

OpenStudy (anonymous):

Let the number of adults b x and the number of children b y. 52x+22y=20690 x+y=500 x=500-y 52(500-y)+22y=20690 26000-52y+22y=20690 5310=30y y=177 x=500-177 =323